OpenStudy (zenmo):

Find the solution of the differential equation that satisfies the given initial condition.

OpenStudy (zenmo):

\[\frac{ dP }{ dt }=5\sqrt{Pt}, P(1) = 2\]

OpenStudy (zenmo):

\[\frac{ 1 }{ \sqrt{P} }\frac{ dP }{ dt }= 5\sqrt{t}*dt\]\[\int\limits_{}^{}P^{-1/2}dP = 5\int\limits_{}^{}t^{1/2}dt\]\[2P^{1/2} = 5[\frac{2 }{ 3 }t^{3/2}] + C\] Let C/2 = K \[P = (\frac{ 5 }{ 3 }t^{3/2}+K)^2 = \frac{ 25 }{ 9 }t^3+\frac{ 10K }{ 3 }t^{3/2}+K^2\] Given condition P(1) = 2, \[\frac{ 25 }{ 9 }(1)^3+\frac{ 10K }{ 3 }(1)^{3/2}+K^2=2 => \frac{ 25 }{ 9 }+\frac{ 30k }{ 9 }+\frac{ 9 }{ 9 }K^2=2\]\[9k^2+30k+25=18 => 9k^2+30k+7 = 0\] \[K = \frac{ -30 \pm \sqrt{900-4(9)(7)} }{ 2(9) } = \frac{ -30 \pm \sqrt{648} }{ 18 }= \frac{ -30 \pm 25 \sqrt{23} }{ 18 }\] \[P = ( \frac{ 5 }{ 3 }t^{3/2}-\frac{ 30 + 25 \sqrt{23} }{ 18 })^2\] My answer is wrong, what do?

Parth (parthkohli):

Why did you square both sides omg\[2\sqrt 2 = \frac{10}3+C\]

OpenStudy (zenmo):

don't you need to square both sides to get rid of P^1/2?

Parth (parthkohli):

You got two values for \(C\) when there is only actually one. Squaring both sides always leads to extraneous solutions.

OpenStudy (inkyvoyd):

how do you know it's the positive one and not the negative one? lol you have a non-linear equation here so you need to make sure you IVP fits one solution and not the other.

OpenStudy (zenmo):

for the positive and negative, I just figure whatever the output condition is: like if P(1) = -2, then it is negative and if P(1) = 2 then it's positive

Parth (parthkohli):

You don't square the sides at all. Just use\[2P^{1/2} = \frac{10}3t^{3/2}\]

Parth (parthkohli):

+ c

OpenStudy (zenmo):

then divide by 2 both sides?

OpenStudy (inkyvoyd):

you can solve for the constant at any time lol... then after you've done so, you can get isolat ethe dependent variable

OpenStudy (zenmo):

\[P^{1/2}=\frac{ 5 }{ 3 }t^{3/2}+\frac{ C }{ 2 }\] Then use the initial condition P(1) = 2 now?

Parth (parthkohli):

Yeah, sure.

OpenStudy (zenmo):

\[\sqrt{2}=\frac{ 5 }{ 3 }+\frac{ C }{ 2 } => \sqrt{2}-\frac{ 5 }{ 3 }= \frac{ C }{ 2 } = > C = 2\sqrt{2}-\frac{ 10 }{ 3 }\]

OpenStudy (zenmo):

\[P = (\frac{ 5 }{ 3 }t^{3/2}+\frac{ 2\sqrt{2}-\frac{ 10 }{ 3 } }{ 2 })^2\] Like that?

Parth (parthkohli):

OK yeah that works

OpenStudy (zenmo):

The website isn't accepting this as an answer, maybe I have to simplify it?

OpenStudy (zenmo):

Source of the question

OpenStudy (astrophysics):

Show's you're multiplying by 2 rather than squaring it

OpenStudy (zenmo):

It's not being multiplied by 2, it is actually raised to the 2nd power. It is difficult to see since the parenthesis are large making the exponent look like a multiplication

OpenStudy (astrophysics):

Maybe try simplifying \[\frac{ 2\sqrt{2}-\frac{ 10 }{ 3 } }{ 2 } = \sqrt{2}-\frac{ 5 }{ 3 }\]

OpenStudy (zenmo):

Yep that was it, had to simplify it.

OpenStudy (astrophysics):

That's silly, but glad it worked out!

OpenStudy (zenmo):

Haha yea, Thanks all!